A three-year-old machine has a cost of $50,000, an estimated residual value of $5,000, and an estimated useful life of five years. The company uses straight-line depreciation. Calculate the net book value at the end of the third year.

a.

A three-year-old machine has a cost of $50,000, an estimated residual value of $5,000, and an estimated useful life of five years. The company uses straight-line depreciation.

Calculate the net book value at the end of the third year.


b.

A three-year-old machine has a cost of $37,600, an estimated residual value of $2,600, and an estimated useful life of 35,000 machine hours. The company uses units-of-production depreciation and ran the machine 5,200 hours in year 1; 6,150 hours in year 2; and 6,800 hours in year 3.

Calculate the net book value at the end of the third year.
Answer #1

a) Annual Depreciation = (50000-5000/5) = 9000

Net book value = Cost-Accumulated Dep = 50000-(9000*3) = 23000


b) Depreciation expense per hour = (37600-2600)/35000 = 1 per hour

Accumulated Dep = (5200+6150+6800)*1 = 18150

Net book value = 37600-18150 = 19450


answered by: Stalicivi